Circular No. 6037

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Using the observed times of the more significant impacts of the come-

tary fragments, S. Nakano, Sumoto, Japan, updates the IAUC 6026 predictions of jovian central-meridian passages of the sites as follows (in order of the first passage on July 25 UT and assuming rotation with system III): site L, July 25.006 UT; G, 25.051; S, 25.059; R, 25.075; Q, 25.093; H, 25.135; E, 25.201; A, 25.236; C, 25.280; K, 25.340; W, 25.346. For the lesser observed impacts he finds: D, July 25.060 UT; B, 25.098; N, 25.107; F, 25.212; M, 25.331. Visual observations (communicated via the SL9 message center) around July 23.9 UT by J. Rogers using the 0.30-m Northumberland refractor at the University of Cambridge Observatories indicate the following: site L, large and dark, in several components; combined G + S (+ D), large and complex; R, tiny black spot south of the main line of spots; Q, single black spot; H, very dark, north of the main line. Summarizing recent reports by other observers, Rogers adds that sites E, A and C are still detectable (with some fading for A) and that the combined K + W site contains three or four dark nuclei (M and supposedly U being also in this region). There do not seem to be confirmed visual sightings of sites B, F (near E), N (near Q), P, Q_{2}, T (near E) or V (near E).

PERIODIC COMET McNAUGHT-HARTLEY (1994n) The following orbital elements, from MPC 23766, confirm this comet

to be of short period (cf. IAUC 6016):

T = 1994 Nov. 11.202 TT  ω = 300.^{o}216
e = 0.61969  Ω = 38.118 }2000.0
q = 2.60969 AU  i = 17.655
a = 6.86205 AU n^{o} = 0.054831 P = 17.98 years
1994 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 27 21^{h}28.^{m}85 −43^{o}50.^{′}2 1.820 2.758 152.^{o}0 10.^{o}0 15.7
Aug. 6 21 22.41 −44 23.8 1.794 2.732 152.1 10.0 15.6
16 21 15.39 −44 32.3 1.792 2.709 148.8 11.2 15.6
26 21 09.01 −44 13.0 1.810 2.688 143.2 13.0 15.6
Sept. 5 21 04.28 −43 26.8 1.849 2.669 136.2 15.2 15.6
15 21 02.00 −42 16.6 1.905 2.653 128.8 17.2 15.6
25 21 02.51 −40 47.3 1.977 2.639 121.2 19.0 15.7
Oct. 5 21 05.79 −39 03.1 2.061 2.628 113.7 20.4 15.8

1994 July 25 Brian G. Marsden